# Talk2Move: Reinforcement Learning for Text-Instructed Object-Level Geometric Transformation in Scenes

### v1

@@ 6.4 Qualitative Evaluation
GPT-Image-1 [openai2024gptimage1] 0.9318 0.4351 0.9453 0.4136 0.9405 0.4278
Flux-Kontext [fluxkontext] 0.9660 0.5704 0.9821 0.5390 0.9739 0.5145
Bagel [bagel] 0.9570 0.3308 0.9356 0.4138 0.9790 0.3027
QwenImageEdit [qwenimage] 0.9728 0.5834 0.9777 0.6000 0.9724 0.5869
@@ 6.5 Ablation Studies
SFT vs RL. Tab.˜3 compares how SFT and RL contribute to spatial manipulation performance. We evaluate both approaches on two backbone models: the widely used editing backbone Flux-Kontext [fluxkontext] and the more advanced QwenImageEdit model [qwenimage], which integrates a VLM with a diffusion generator. To ensure a fair comparison, all experiments are trained solely on object translation data.
@@
Figure 5: Qualitative results on object translation, rotation and resize over state-of-the-art image editing models. For each task, we provide one real image editing result (source from OpenImagesV6 [OpenImages2]) and one synthetic image editing result to showcase the generalization ability of Talk2Move.
@@
Ablation on Reward Model. We conduct an ablation comparing a VLM-based reward model with our spatial reward derived from a specialist model (Orient-Anything [orientanything] + rule-based evaluation). Editing quality is assessed using Qwen2.5-VL-Instruct [Qwen2.5-VL] to measure alignment between the edited image and the prompt. As shown in Fig.˜4(c), both reward models improve evaluation reward during GRPO training. However, the VLM-based reward often produces overly optimistic and unstable scores, yielding a rotation error of 0.3294 and an accuracy of 11.63% according to evaluation. In contrast, our spatial reward leads to a lower error of 0.2861 and a much higher accuracy of 29.55%, indicating that it provides a more reliable supervision signal for GRPO-based spatial manipulation.

### v2

@@ 6.4 Qualitative Evaluation
GPT-Image-1 [26] 0.9318 0.4351 0.9453 0.4136 0.9405 0.4278
Flux-Kontext [12] 0.9660 0.5704 0.9821 0.5390 0.9739 0.5145
Bagel [4] 0.9570 0.3308 0.9356 0.4138 0.9790 0.3027
QwenImageEdit [39] 0.9728 0.5834 0.9777 0.6000 0.9724 0.5869
@@ 6.5 Ablation Studies
SFT vs RL. Tab.˜3 compares how SFT and RL contribute to spatial manipulation performance. We evaluate both approaches on two backbone models: the widely used editing backbone Flux-Kontext [12] and the more advanced QwenImageEdit model [39], which integrates a VLM with a diffusion generator. To ensure a fair comparison, all experiments are trained solely on object translation data.
@@
Figure 5: Qualitative results on object translation, rotation and resize over state-of-the-art image editing models. For each task, we provide one real image editing result (source from OpenImagesV6 [11]) and one synthetic image editing result to showcase the generalization ability of Talk2Move.
@@
Ablation on Reward Model. We conduct an ablation comparing a VLM-based reward model with our spatial reward derived from a specialist model (Orient-Anything [38] + rule-based evaluation). Editing quality is assessed using Qwen2.5-VL-Instruct [1] to measure alignment between the edited image and the prompt. As shown in Fig.˜4(c), both reward models improve evaluation reward during GRPO training. However, the VLM-based reward often produces overly optimistic and unstable scores, yielding a rotation error of 0.3294 and an accuracy of 11.63% according to evaluation. In contrast, our spatial reward leads to a lower error of 0.2861 and a much higher accuracy of 29.55%, indicating that it provides a more reliable supervision signal for GRPO-based spatial manipulation.